If one GM 'g' and two AM's 'p' and 'q' are inserted between two numbers a and b, then prove that :-


(2p-q)(p-2q)= -g^2

Grade:

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear sridhar ,

g = (ab)1/2  is quite simple.

now a , p , q, b are in AP .

b =  a +(4-1) d  ;   d = b -a/3

thereofore ,  p =  a +(b-a/3) =  2a+b/3

                     q  = p +d =  (2a+b/3) + (b -a)/3 = 2b+a/3

so 2p-q  = 3a/3 =a  and p -2q  = -b

so , (2p-q)(p-2q) =  -ab =  -g^2

proved . this is the straight forward  method and i think is most intuitive so i suggest follow this method.

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

 

Win exciting gifts by answering the questions on Discussion Forum.

 

                                                                    
                                                                    
                                                                    
                                            
 






 

 

If one GM 'g' and two AM's 'p' and 'q' are inserted between two numbers a and b, then prove that :-

(2p-q)(p-2q)= -g^2